OpenStudy (anonymous):

Can someone help me turn these equations into slope intercept form? y=6x-5 y=3x-2 x=3x+2 x=2x-1 x+y=1 4x+4y=4

OpenStudy (zzr0ck3r):

solve for y, and you are done

OpenStudy (johnweldon1993):

Well a way to look at it is...the first 2 already are in that form...so try and make the others look like that

OpenStudy (zzr0ck3r):

the third and fourth don't make sense for the question.

OpenStudy (anonymous):

I don't understand how to do the others because some of them don't have y's

OpenStudy (zzr0ck3r):

unless you consider x = a point slope form

OpenStudy (johnweldon1993):

If they don't have y's...they cannot be solved for y and therefor cannot be turned into slope intercept form

OpenStudy (anonymous):

> I think I finally figured out the last part of assignment 6.01. > > Intersecting lines- y=6x-5 > y=3x-2 > Solution= (1,1) > These intersect at only one point. The system of linear equations only have one solution. > > Parallel lines- x=3x+2 > x=2x-1 > Solution= No solution > These lines never intersect because they are parallel lines. They have no points in common. > > The same line- x+y=1 > 4x+4y=4 > Solution=Infinity > These lines are exactly the same. They lay directly on top of each other. This was my homework but then my teacher told me to turn the equations into slope intercept form.

OpenStudy (zzr0ck3r):

correct, they would have "infinite" slope, so this does not work

OpenStudy (anonymous):

so for 3 and 4 I don't have to turn them into slope intercept form?

OpenStudy (johnweldon1993):

You cannot ...no, infinite slope = no slope intercept form

OpenStudy (anonymous):

what about the 5th and 6th one?

OpenStudy (johnweldon1993):

x + y = 1 how would you solve for 'y' ...?

OpenStudy (johnweldon1993):

Notice how in the first and second one....it is just y = ...something *the 'y' is by itself on that left side* x + y = 1 we need to get 'y' alone on that left side....

OpenStudy (anonymous):

Do I minus the x from one?

OpenStudy (johnweldon1993):

That would be correct....you would subtract 'x' from both sides x + y = 1 -x -x y = -x + 1 this is your slope intercept form

OpenStudy (anonymous):

Yes! thanks.

OpenStudy (johnweldon1993):

And for the 6th one...?

OpenStudy (anonymous):

minus 4x from 4?

OpenStudy (johnweldon1993):

again correct...subtract 4x from both sides 4x + 4y = 4 -4x -4x 4y = -4x + 4 now we dont want 4y....we want 'y'...so what do we do now...?

OpenStudy (anonymous):

Minus 4y from 4?

OpenStudy (johnweldon1993):

not quite...to get 4y to equal y....we want to divide it by 4 right? because \[\frac{ 4y }{ 4 }\] this equals 1y right? 4 only goes into 4 once.... so we divide everything by 4 4y = -4x + 4 ------------- 4 4 4 y = ......?

OpenStudy (anonymous):

y=-1x+1?

OpenStudy (johnweldon1993):

correct....just remember you really don't need to put -1x y = -x + 1 would be the same thing...regardless...you are correct

OpenStudy (anonymous):

Thanks so much! I suck at algebra. I prefer geometry. lol

OpenStudy (johnweldon1993):

Wow, we are opposite...I despise geometry...and algebra is great lol

OpenStudy (anonymous):

Yeah it's usually that way with people. Geometry is more memory and algebra is all work and numbers.